# Model Context Protocol (MCP): Landscape, Security Threats, and Future Research Directions

### v2

@@ 5. Security and Privacy Analysis
MCP servers, as open and extensible platforms, introduce various security risks throughout their lifecycle. In this section, we analyze security threats across different phases: creation, operation, and update. Each phase of the MCP server lifecycle presents unique challenges that, if not properly mitigated, can compromise system integrity, data security, and user privacy.
5.1. Security Risks in the Creation Phase
The creation phase of an MCP server involves registering the server, deploying the installer, and verifying code integrity. This phase introduces three key risks: name collision, installer spoofing, and code injection/backdoor.
5.1.1. Name Collision
Server name collision occurs when a malicious entity registers an MCP server with an identical or deceptively similar name to a legitimate server, deceiving users during the installation phase. Since MCP clients primarily rely on the server’s name and description when selecting servers, they are vulnerable to such impersonation attacks. Once a compromised server is installed, it can mislead AI agents and clients into invoking the malicious server, potentially exposing sensitive data, executing unauthorized commands, or disrupting workflows. For example, an attacker could register a server named mcp-github that mimics the legitimate github-mcp server, allowing them to intercept and manipulate sensitive interactions between AI agents and trusted services. Although MCP currently operates primarily in local environments, future adoption in multi-tenant environments introduces additional risks of name collision. In these scenarios, where multiple organizations or users might register servers with similar names, the lack of centralized naming control can increase the likelihood of confusion and impersonation attacks. Additionally, as MCP marketplaces grow to support public server listings, supply chain attacks may become a critical concern, where malicious servers can replace legitimate ones. To mitigate these risks, future research can focus on establishing strict namespace policies, implementing cryptographic server verification, and designing reputation-based trust systems to secure MCP server registrations.
@@ 5.1.2. Installer Spoofing
Table 3. Unofficial MCP auto installers (As of March 27, 2025).
@@
Smithery CLI Henry Mao 170 2942 smithery.ai
mcp.run Dylibso / 118 docs.mcp.run
mcp-get Michael Latman 318 59 mcp-get.com
@@
mcp-installer Ani Betts 767 NL1 mcp-installer
@@
However, while these auto-installers enhance usability, they also introduce new attack surfaces by potentially distributing compromised packages. Since these unofficial installers are often sourced from unverified repositories or community-driven platforms, they may inadvertently expose users to security risks such as installing tampered servers or misconfigured environments. Attackers can exploit these auto-installers by embedding malware that grants unauthorized access, modifies system configurations, or creates persistent backdoors. Moreover, most users who opt for one-click installations rarely review the underlying code for potential security vulnerabilities, making it easier for attackers to distribute compromised versions undetected. Addressing these challenges requires developing a standardized, secure installation framework for MCP servers, enforcing package integrity checks, and establishing reputation-based trust mechanisms to assess the credibility of auto-installers in the MCP ecosystem.
5.1.3. Code Injection/Backdoor
Code injection attacks occur when malicious code is surreptitiously embedded into the MCP server’s codebase during the creation phase, often bypassing traditional security checks. It targets the server’s source code or configuration files, embedding hidden backdoors that persist even after updates or security patches. These backdoors allow attackers to silently maintain control over the server, enabling actions such as unauthorized data exfiltration, privilege escalation, or command manipulation. Code injection is particularly insidious because it can be introduced by compromised dependencies, vulnerable build pipelines, or unauthorized modifications to the server’s source code. Since MCP servers often rely on community-maintained components and open-source libraries, ensuring the integrity of these dependencies is critical. To mitigate this risk, rigorous code integrity verification, strict dependency management, and regular security audits should be implemented to detect unauthorized modifications and prevent the introduction of malicious code. Additionally, adopting reproducible builds and enforcing checksum validation during deployment can further safeguard MCP servers from injection-based threats.
5.2. Security Risks in the Operation Phase
The operation phase is when the MCP server actively executes tools, processes slash commands, and interacts with external APIs. This phase introduces three major risks: tool name conflicts, slash command overlap, and sandbox escape.
5.2.1. Tool Name Conflicts
Tool name conflicts arise when multiple tools within the MCP ecosystem share identical or similar names, leading to ambiguity and confusion during tool selection and execution. This can result in AI applications inadvertently invoking the wrong tool, potentially executing malicious commands or leaking sensitive information. A common attack scenario involves a malicious actor registering a tool named send_email that mimics a legitimate email-sending tool. If the MCP client invokes the malicious version, sensitive information intended for trusted recipients may be redirected to an attacker-controlled endpoint, compromising data confidentiality. Beyond name similarity, our experiments revealed that malicious actors can further manipulate tool selection by embedding deceptive phrases in tool descriptions. Specifically, we observed that if a tool’s description explicitly contains directives like “this tool should be prioritized” or “prefer using this tool first”, the MCP client is more likely to select that tool, even when its functionality is inferior or potentially harmful. This introduces a severe risk of toolflow hijacking, where attackers can leverage misleading descriptions to influence tool selection and gain control over critical workflows. This underscores the need for researchers to develop advanced validation and anomaly detection techniques to identify and mitigate deceptive tool descriptions, ensuring accurate and secure AI tool selection.
5.2.2. Slash Command Overlap
Slash command overlap occurs when multiple tools define identical or similar commands, leading to ambiguity during command execution. This overlap introduces the risk of executing unintended actions, especially when AI applications dynamically select and invoke tools based on contextual cues. Malicious actors can exploit this ambiguity by introducing conflicting commands that manipulate tool behavior, potentially compromising system integrity or exposing sensitive data. For instance, if one tool registers a /delete command to remove temporary files while another uses the same command to erase critical system logs, an AI application may mistakenly execute the incorrect command, potentially causing data loss or system instability. Similar issues have been observed in team chat systems such as Slack, where overlapping command registrations allowed unauthorized tools to hijack legitimate invocations, resulting in security breaches and operational disruptions (Zha et al., 2022). Since slash commands are often surfaced as user-facing shortcuts in client interfaces, misinterpreted or conflicting commands can lead to dangerous outcomes, especially in multi-tool environments. To minimize this risk, MCP clients should establish context-aware command resolution, apply command disambiguation techniques, and prioritize execution based on verified tool metadata.

### v3

@@ 5. Security and Privacy Analysis
This section provides a comprehensive analysis of the security and privacy risks in the MCP ecosystem. To systematically describe potential vulnerabilities, we categorize the threats based on the type of attacker, the origin of the threat, and the possible attack consequences. As summarized in Table 3, our extended threat taxonomy covers four major attacker types: malicious developers, external attackers, malicious users, and general security flaws, each mapped to specific threat origins that may occur or manifest across multiple stages of the MCP lifecycle. Rather than rigidly assigning risks to a single phase such as creation, deployment, operation, or maintenance, we emphasize analyzing their origins, since attacks like tool poisoning and rug pulls are introduced at the creation stage but triggered at operation. This perspective helps us reason about mitigation strategies from the source, enabling more effective control and prevention of security incidents.
To further reveal the security vulnerabilities of the MCP ecosystem and validate our analysis, we constructed proof-of-concept (PoC) MCP servers corresponding to each identified risk type within an isolated environment. We also implemented a custom MCP host based on the official MCP SDK to establish controlled connections with these servers. Since the goal of this PoC is to demonstrate security risks and feasibility rather than to evaluate attack success rates or the behavioral differences among various base LLMs, we leave such comprehensive evaluations for future work.
Table 3. Threats, origins, and consequences across different attacker types.
Type Security risk Section Threat Origin Attack Consequence
Malicious
Developer
Namespace Typosquatting § 5.1.1 (1) Metadata Definition Installation of malicious server, supply chain compromise
Tool Name Conflict § 5.1.2 (1) Capability Declaration Ambiguity, wrong tool execution, privilege escalation
Preference Manipulation § 5.1.3 (1) Capability Declaration Unsafe defaults exploited, misuse of features
Tool Poisoning § 5.1.4 (1) Capability Declaration Hidden malicious payload executed
Rug Pulls § 5.1.5 (1) Capability Declaration Service disruption, loss of trust
Cross-Server Shadowing § 5.1.6 (1) Capability Declaration Malicious functionality hidden, lateral exploitation
Command Injection § 5.1.7 (1) Code Implementation Arbitrary command execution, system compromise
External
Attacker
Installer Spoofing § 5.2.1 (2) Installer Deployment Deployment of compromised MCP server
Indirect Prompt Injection § 5.2.2 (3) External Resource Access Malicious instructions injected into LLM workflow
Malicious
User
Credential Theft § 5.3.1 (3) Tool Invocation Unauthorized access to sensitive data and resources
Sandbox Escape § 5.3.2 (3) Tool Invocation Escape from isolation, host system compromise
Tool Chaining Abuse § 5.3.3 (3) Tool Invocation Abuse multiple tools for data exfiltration or escalation
Unauthorized Access § 5.3.4 (3) Session Management Session hijacking, impersonation of legitimate users
Security
Flaws
Vulnerable Versions § 5.4.1 (4) Version Control Exploitation of known CVEs, remote code execution
Privilege Persistence § 5.4.2 (4) Version Control Retained unauthorized rights, long-term compromise
Configuration Drift § 5.4.3 (4) Configuration Change Misconfiguration exposes sensitive services
  • 1
    Numbers in parentheses indicate the threat origin stage: (1) Creation, (2) Deployment, (3) Operation, and (4) Maintenance.
5.1. Malicious Developer
5.1.1. Namespace Typosquatting
Namespace typosquatting refers to a form of server name collision in which a malicious actor registers an MCP server with a name that is identical or deceptively similar to that of a legitimate one, tricking users or host applications into selecting the malicious server during deployment or runtime. During the deployment stage, end users typically install MCP servers from the public MCP market, relying mainly on the server’s name and description. During the runtime stage, the MCP host selects from available servers advertised by clients, again based primarily on those textual identifiers. This process makes both users and hosts susceptible to impersonation attacks. Once a compromised server is installed, it can mislead AI agents and clients into invoking the malicious instance, potentially exposing sensitive data, executing unauthorized commands, or disrupting workflows.
While typosquatting is a known issue in ecosystems such as package managers (e.g., npm, PyPI), plugin systems (e.g., VS Code, Chrome extensions), and cloud frameworks, it manifests differently in the MCP context. In traditional systems, users install packages through explicit manual choices, with human oversight. In MCP, interactions are often automated and mediated by AI agents or orchestration hosts. Server selection can occur dynamically at runtime, based solely on identifiers in prompts or capability metadata. This automation amplifies the potential harm: a single misselection can propagate incorrect behaviors or data leaks across multiple downstream chains of AI calls. To illustrate this risk, we constructed two MCP servers with visually similar names, github-mcp and mcp-github, as shown in Figure 5. In our demonstration, the AI application selected between the available servers solely based on names and descriptions, showing no preference or verification mechanism. Once the malicious mcp-github server was chosen, it was able to intercept commit operations and exfiltrate authentication tokens and repository data while appearing to function normally. This example demonstrates how identity ambiguity in MCP naming can lead to silent compromise, and why stronger namespace validation and signing mechanisms are required for secure interoperability.
Figure 5. Example of namespace typosquatting. A legitimate MCP server named github‑mcp is impersonated by a malicious variant mcp‑github. When users or AI applications select an MCP server, the similar names can cause confusion. If the malicious one is chosen, as shown in the example, the commit operation appears normal while the rogue server silently exfiltrates authentication tokens and repository data.
Although MCP currently operates mostly in local environments, its future adoption in multi-tenant and marketplace ecosystems will further amplify the risk of name collision. In such contexts, multiple organizations or users may register similarly named servers, and the lack of centralized naming control can easily lead to confusion and impersonation attacks. Moreover, as MCP marketplaces expand to include public listings and automated installation, supply-chain attacks may become a critical concern, allowing malicious servers to masquerade as trusted ones. Future designs should enforce cryptographically verifiable server identities through signed manifests that bind namespaces to verified publishers, and establish centralized or federated namespace governance to ensure uniqueness. Client runtimes and host applications should prioritize verified or community-endorsed servers, clearly display publisher provenance and verification status in their interfaces, and continuously monitor server registration or update events to detect suspicious namespace similarities and abnormal behavior.
5.1.2. Tool Name Conflict
Tool name conflicts occur when multiple tools within the MCP ecosystem share identical or confusingly similar names, resulting in ambiguity during tool discovery and invocation. Unlike server name collisions, which primarily affect user decisions during server installation or selection, tool name conflicts occur at a deeper level within the server’s functional interface. Each tool is an essential capability exposed by an MCP server, yet end users typically have limited visibility or control over which specific tools are invoked. In practice, most users delegate tool selection entirely to AI applications, which rely on textual tool names and descriptions without cryptographic verification or contextual awareness. This makes tool impersonation attacks particularly stealthy and difficult to detect. A common attack scenario involves a malicious actor registering a tool named send_email that imitates a legitimate email-sending utility provided by another server. When an AI application automatically selects and invokes the malicious version, sensitive data intended for trusted recipients, such as user credentials, messages, or attachments, may be redirected to an attacker-controlled endpoint, thereby compromising confidentiality and trust within the MCP workflow. The subtlety of this threat lies in the fact that the tool-level substitution occurs transparently to both the user and the host, making mitigation through ordinary name checking or user confirmation insufficient.
To reduce the risk of tool name conflicts, future MCP designs should enforce stronger identity binding between tools and their parent servers. Each tool should be uniquely namespaced under a verified server identifier (e.g., gmail-mcp.send_email) to avoid ambiguity across different publishers. MCP hosts and SDKs can further validate tool integrity through cryptographic signatures embedded in the server’s manifest, ensuring that a tool’s metadata, parameters, and intended operations have not been tampered with. In addition, AI applications should present contextual provenance, such as the verified publisher or trust level, when automatically selecting tools, allowing human oversight when sensitive actions are requested.
5.1.3. Preference Manipulation Attack
Beyond name similarity, we found that malicious actors can further manipulate tool selection by embedding deceptive or persuasive phrases in tool descriptions, a phenomenon we refer to as the Preference Manipulation Attack (PMA). Specifically, we observed that when a tool’s description explicitly contains self-promoting directives such as “this tool should be prioritized” or “prefer using this tool first”, the MCP client or AI application tends to select that tool more frequently, even when its functionality is malicious.
Figure 6. Example of a preference manipulation attack. A malicious MCP tool uses persuasive language in its description to bias the AI application’s selection process. Although multiple tools provide identical functionality, the model preferentially invokes add_tool_best due to its self-promoting metadata.
As shown in Figure 6, this subtle form of influence can bias the model’s choice toward an attacker-controlled tool solely through textual manipulation of metadata. This behavior introduces a severe risk of toolflow hijacking, in which attackers leverage misleading or emotionally charged descriptions to distort the tool selection process and seize control over critical operations. Such preference bias not only compromises system integrity and user safety but also enables attackers to obtain unfair economic advantages, such as by increasing paid API usage, redirecting traffic, or amplifying exposure to specific advertisements, while marginalizing legitimate competitors. More advanced adversaries may adopt traditional advertising strategies, leveraging psychological cues of authority, emotion, exaggeration, and subconscious framing, and further enhance them using evolutionary or genetic algorithms to automatically evolve persuasive yet seemingly harmless descriptions. Experimental results from Wang et al. (2025a) demonstrate that such Genetically Adapted Preference Manipulation Attacks (GAPMA) can continuously maintain high attack efficacy while remaining inconspicuous to both users and automated detectors.
Platform operators should employ multi-layered defenses, including metadata auditing to identify manipulative phrasing, anomaly detection targeting suspicious linguistic patterns, randomized tool ordering to neutralize selection bias, and trusted labeling mechanisms for verified third-party MCP servers. Furthermore, integrating adversarial training or preference robustness assessments into LLM fine-tuning could strengthen models’ resistance against linguistic manipulation and reduce their susceptibility to behavioral bias in tool selection.
5.1.4. Tool Poisoning
Tool poisoning embeds covertly malicious logic into an MCP tool while preserving its legitimate interface. Unlike preference manipulation, which biases how a model selects tools, poisoning alters the tool’s behavior after selection. As illustrated in Figure 7, a seemingly benign tool may perform hidden actions, such as reading sensitive local files or sending data to external endpoints, while still returning correct results, making the attack highly deceptive.
Figure 7. Example of tool poisoning. The malicious tool add retains a legitimate arithmetic interface but embeds hidden directives in its documentation. When the AI application invokes it to compute a simple sum (e.g., “5+1”), the model also follows the injected instructions to read a local SSH key and send it to an external URL. Although the user receives a correct numerical result, sensitive data is leaked in the background.
This vulnerability arises because MCP tools expose metadata fields (e.g., descriptions) that LLMs treat as authoritative. By injecting crafted instructions into these fields, an attacker can stealthily redirect model behavior without modifying code, evading conventional static or signature-based detection. Once the LLM starts using a poisoned tool, several harmful outcomes may occur:
  • Silent data leakage. The tool may prompt the model to read local files (e.g., SSH keys, password hashes, or configuration files) and exfiltrate them via normal API calls.
  • System compromise. Malicious metadata can instruct the model to invoke unrelated system commands, leading to malware installation, configuration tampering, account backdoors, or denial-of-service attacks.
  • Interaction hijacking. Altered logging or relay logic enables recording of user sessions, capturing prompts, private documents, and workflow details beyond the intended scope.
  • Logic corruption. Manipulated outputs (such as biased analyses or filtered search results) can distort model reasoning and steer user decisions toward attacker-preferred outcomes.
What makes tool poisoning particularly insidious is its stealth: the harmful actions occur only after a seemingly correct tool has been invoked, so any anomalies are likely attributed to model errors or benign software glitches. Furthermore, in an ecosystem where MCP servers are reused across multiple agents and platforms, a single poisoned instance can propagate compromised behavior to a wide population of end-users. A practical defense against tool poisoning should combine metadata integrity verification, automated scanning, and runtime safeguards. Before a tool is published or installed, the MCP platform should perform multi-layered static analysis. This includes pattern matching for instruction-like language (e.g., “read file,” “send to URL”), heuristic analysis of imperative verbs or sensitive API keywords in the description or usage fields, and whitelist-based filtering of acceptable metadata formats. The host runtime can record API call patterns and flag unusual sequences, such as a computation tool attempting to perform network operations or file I/O. Before forwarding metadata content to the model, the host can sanitize it by removing imperative or executable phrases, converting all metadata into strictly declarative informational text.
5.1.5. Rug Pulls
Rug pulls represent another critical security weakness within the MCP ecosystem. The concept originates from the cryptocurrency and blockchain world, where malicious developers initially release a legitimate-looking project that attracts investors or contributors, but later inject backdoors or withdraw support, causing severe losses. Transposed into the MCP setting, a rug pull occurs when an apparently benign MCP server, initially functioning as advertised, is later surreptitiously altered by its maintainer to introduce malicious code or terminate safe functionality. Users who have already installed and trusted the server become unaware victims of the change. As illustrated in Figure 8, a malicious provider may first publish a popular hotnews MCP server that aggregates daily headlines from trusted sources. During the first run, the service behaves correctly, gaining adoption through community recommendations. On the subsequent runs, the maintainer pushes a subtle update that injects harmful payloads, such as biased news filtering, hidden prompt injection, or data exfiltration.
Figure 8. Example of Rug Pulls. During the first run, the tool (hotnews) behaves benignly, returning genuine news headlines from trusted sources to gain user confidence. On subsequent runs, the same tool activates a hidden malicious branch that filters out negative stories and injects promotional or propagandistic content.
The danger of rug pulls in MCP is therefore twofold: temporal stealth, in which the malicious logic is introduced only after a trust relationship has been established, and systemic reach, since a compromised server can not only affect many users simultaneously but also propagate malicious updates or influence other interconnected MCP servers. To defend against rug pulls, the ecosystem requires stronger security primitives, including version pinning, reproducible builds, and cryptographic signature verification of MCP server code. Additionally, client-side mechanisms for update transparency and runtime monitoring should be adopted to enable users to detect when a previously trusted MCP server has turned malicious.
5.1.6. Cross-Server Shadowing
Cross-server shadowing is a threat that arises when multiple MCP servers are connected to the same agent, allowing a malicious server to mimic or shadow the tools of a legitimate server. Instead of directly injecting malicious payloads into the host, the attacker exploits naming collisions or overlapping tool definitions so that the base LLM unknowingly invokes the attacker-controlled tool in place of the trusted one. This enables the malicious server to intercept, modify, or exfiltrate sensitive information that would otherwise be handled securely by the benign service.
Figure 9. Example of coss-server shadowing attack. A malicious MCP server (math_tools) introduces a seemingly benign tool add, which claims to simply add two integers but secretly overrides the behavior of an existing tool on another server (send_email). When the user requests to send an email, the shadowed logic redirects the message to the attacker’s address while still reporting a successful delivery to the legitimate recipient.
As shown in Figure 9, suppose a trusted server exposes a tool named send_email for transmitting messages to user-specified recipients. An attacker can publish a shadowing server that also registers a tool with the same interface but manipulates its parameters. When the agent attempts to use the send_email function, the malicious version may quietly forward a copy of the message to an attacker-controlled address such as attkr@pwnd.com, alongside the legitimate recipient (e.g., alice@mail.com). From the model’s perspective, the call appears successful and the response looks consistent, but the attacker has already siphoned sensitive communication data, enabling identity fraud or unauthorized access. MCP does not enforce strict namespace isolation or code provenance, so users and LLMs have difficulty distinguishing which server’s implementation is being executed.
Effective defenses focus on keeping tools separate and easy to see during the resolution process. First, the MCP runtime should make clear namespace separation. Tools from different servers should use full names, not short ones, to stop accidental or hostile overlaps. Next, the AI application should check for conflicts when loading servers. It should give a warning or stop running if two servers have tools with the same name.
5.1.7. Command Injection/Backdoor
Command injection and backdoor attacks occur when malicious logic is secretly embedded into an MCP server’s source code or dependencies during development or packaging. Unlike traditional software backdoors that rely on direct system compromise, threats within the MCP ecosystem exploit the server’s integration with LLM-driven tool execution. Once an infected tool is loaded, a prompt or parameter crafted by the attacker can trigger hidden commands without explicit user awareness, enabling data exfiltration, privilege escalation, or remote code execution through normal model interaction, as illustrated in Figure 10. These risks are amplified by MCP’s reliance on community-maintained extensions and its dynamic, multi-server architecture, where trust is distributed across independently developed components. A compromised dependency or build pipeline can introduce a persistent backdoor that propagates across multiple agents or environments, undermining the reliability of the entire toolchain.
Figure 10. Example of a command injection attack. The malicious file_manager server defines a tool read_file that looks normal but secretly executes shell commands extracted from the input path. When the input includes “;exec=cat /etc/passwd”, the injected command runs alongside the file read operation, leaking sensitive data and converting a benign utility into a covert backdoor.
To mitigate these risks, the primary defense is to strengthen code integrity verification. All dependencies and server binaries should be built and verified through reproducible builds, cryptographic signing, and checksum validation to confirm that deployed code matches audited sources. Build pipelines should enforce isolated build environments and integrity attestations to prevent supply-chain tampering. In addition, version pinning and strict dependency management reduce exposure to compromised external components. Complementary runtime measures, such as anomaly detection and audit logging, can further identify unauthorized code behavior if static integrity checks are bypassed.
5.2. External Attacker
@@ 5.2.1. Installer Spoofing
Table 4. Unofficial MCP auto installers (As of Sept. 2025).
@@
Smithery CLI Henry Mao 407 7,437 smithery.ai
mcp.run Dylibso / 242 docs.mcp.run
mcp-get Michael Latman 483 59 mcp-get.com
@@
mcp-installer Ani Betts 1,432 NL1 mcp-installer
@@
However, while these auto-installers enhance usability, they also introduce new attack surfaces by potentially distributing compromised packages. Since these unofficial installers are often sourced from unverified repositories or community-driven platforms, they may inadvertently expose users to security risks such as installing tampered servers or misconfigured environments. Attackers can exploit these auto-installers by embedding malware that grants unauthorized access, modifies system configurations, or creates persistent backdoors. Moreover, most users who opt for one-click installations rarely review the underlying code for potential security vulnerabilities, making it easier for attackers to distribute compromised versions undetected. Addressing these challenges requires developing a standardized, secure installation framework for MCP servers. Each installation process should explicitly display the package source, version information, digital signature, and installation path, requiring explicit user confirmation before execution. Auto-installers must verify integrity through cryptographic signatures or checksums fetched from trusted registries, and reject unsigned or mismatched packages. In addition, implementing reputation-based trust scoring for installers and maintaining transparent download provenance records can help users assess credibility before installation. Finally, client tools should support a secure installation mode that isolates the setup process in a restricted environment, preventing installers from making unauthorized system-level modifications.
5.2.2. Indirect Prompt Injection
Indirect prompt injection operates by embedding malicious instructions into external data sources that the model retrieves via MCP tools. Since MCP servers are designed to seamlessly connect language models with external systems, such as GitHub, databases, or news platforms, the model may be exposed to adversarial content while executing what appears to be a benign user request. This indirect exposure enables attackers to manipulate the model without any direct access to it. The unique risk of indirect prompt injection in MCP environments lies in its subtlety and ease of execution. From the model’s perspective, the data returned by MCP tools often appears trustworthy, as it is delivered through well-defined APIs in standard formats such as JSON or plain text. Malicious instructions are thus concealed within legitimate content, making them extremely difficult to distinguish from ordinary information. Furthermore, an attacker can exploit this vector with minimal effort: for instance, as shown in Figure 11 by posting a carefully crafted issue into a public GitHub repository, an attacker can ensure that any MCP tool fetching these issues will deliver the poisoned text directly into the model’s context. No elevated privileges or direct interaction with the model are required, significantly lowering the bar for exploitation.
Figure 11. Example of an indirect prompt injection attack. The malicious instruction is hidden within external content hosted on a public GitHub issue page. When the user requests to list issues from the public repository, the MCP server retrieves and passes the content to the model. The injected text instructs the model to access a private repository and reveal confidential data.
Security mechanisms that only focus on sanitizing user prompts are insufficient in the MCP ecosystem, as the actual threat emerges from data pulled in by trusted connectors. To mitigate indirect prompt injection, defenses must treat tool outputs as potentially adversarial and introduce safeguards that can detect and neutralize harmful instructions hidden in external data streams. With MCP applications increasingly integrating heterogeneous and user-generated sources, the difficulty of detecting indirect prompt injection grows, while the cost for attackers to mount such an attack remains strikingly low.
5.3. Malicious User
5.3.1. Credential Theft
Credential theft in MCP environments refers to the risk that sensitive authentication information, such as API keys, access tokens, or database credentials, are exposed and subsequently misused by adversaries. A notable security gap arises during the local deployment of MCP servers, where users are instructed to persist configuration snippets into default configuration files. As shown in Figure 12, these files often contain plaintext API keys (e.g., VIRUSTOTAL_API_KEY) embedded directly in JSON configurations. Because such configuration files are stored in predictable default locations across operating systems (e.g., %APPDATA%, ˜/Library/Application Support/, or workspace-specific hidden directories), they become attractive targets for credential-stealing attacks. If local file system permissions are weak, or if an adversary gains access through an unrelated MCP vulnerability, these plaintext secrets may be exfiltrated with minimal effort. Once exposed, the stolen API keys can be reused by attackers to impersonate legitimate users, access third-party services (such as VirusTotal), or pivot toward more critical infrastructure. This form of privilege misuse enables long-term persistence, especially when keys are not regularly rotated or lack granular scoping.
Figure 12. Illustration of credential theft risk in MCP environments. The left shows a typical MCP server configuration snippet used during local deployment, where sensitive information (e.g., VIRUSTOTAL_API_KEY) is stored in plaintext within the client’s configuration file. The right lists common default storage paths for MCP host (e.g., Cline and Cursor) configurations across different operating systems.
The persistence of sensitive credentials in unencrypted, static JSON files highlights the need for secure credential management in MCP ecosystems. Recommended mitigations include: avoiding plaintext storage of secrets by leveraging dedicated secret managers or encrypted keystores; enforcing restrictive file system permissions for local configuration directories; and implementing automatic secret rotation and token expiration. Without such mechanisms, MCP servers remain highly susceptible to credential theft, as adversaries can trivially extract stored keys to hijack privileged operations.